Two identical satellites orbit the earth in stable orbits. One satellite orbits with a speed v at a distance r from the center of the earth. The second satellite travels at a speed that is less than v.At what distance from the center of the earth does the second satellite orbit?


At a distance that is less than r.

At a distance equal to r.

At a distance greater than r.

Now assume that a satellite of mass m is orbiting the earth at a distance r from the center of the earth with speed ve. An identical satellite is orbiting the moon at thesame distance with a speed vm. How does the time Tm it takes the satellite circling the moon to make onerevolution compare to the time Te it takes the satellite orbiting the earth to make onerevolution?

Tm is less than Te.  

Tis equal to Te

Tis greater than Te.
Answer #1
Concepts and reason

The concepts used to solve this problem are orbital speed, and satellite’s time period of revolution.

Initially, use the orbital speed expression to determine the condition of second satellite’s orbit.

Finally, use the time period expression to determine the time period of the satellite.

Fundamentals

Orbital speed of a satellite is the speed with which it goes around the earth in a periodic motion. The path of this motion is called as orbit.

Expression for the orbital speed is,

v=GMrv = \sqrt {\frac{{GM}}{r}}

Here, orbital speed is vv, universal gravitational constant is GG, mass of earth is MM, and the radius of orbit of the satellite is rr.

The time period of satellite’s orbit is the time taken by the satellite to complete one orbit around the planet.

Expression for time period is,

T=2πrvT = \frac{{2\pi r}}{v}

Here, time period of orbit is TT, radius of orbit is rr, and orbital speed is vv.

(1)

The two satellites have equal masses. One satellite is orbiting with an orbital speed vv, and radius rr. The second satellite has an orbital speed vv' that is less than vv.

The expression for orbital speed of the first satellite is,

v=GMrv = \sqrt {\frac{{GM}}{r}}

The condition for the orbital speed of the second satellite is,

v<vv' < v …… (1)

From the expression for the orbital speed,

vα1rv\alpha \frac{1}{{\sqrt r }} …… (2)

The orbital speed is inversely proportional to the square root of the radius of the orbit.

From the expressions (1) and (2), if the orbital speed of second satellite is less than the orbital speed of the first satellite, then the radius of the second satellite should be greater than that of the first satellite.

The incorrect options are,

• At a distance that is less than rr

• At a distance equal to rr

Use the conditionv<vv' < v,

• At a distance that is less than rr

This option is incorrect because the velocity of second satellite is less than the first satellite velocity and hence its radius of orbit cannot be less than that of first satellite’s radius of orbit.

• At a distance equal to rr

This option is incorrect as the orbital speed of both satellites are different and hence they cannot have same radius of orbit.

Hence, the correct option is

• At a distance greater than rr

The above option is correct because, the radius of orbit is inversely proportional to the orbital speed and hence its radius will be greater than that of first satellite.

(2)

Satellite of mass mm goes around the earth in an orbit having radius rr and orbital speed ve{v_e}.

The expression for the orbital speed for this orbit is,

ve=GMer{v_e} = \sqrt {\frac{{G{M_e}}}{r}} …… (3)

If the satellite goes around moon in an orbit having same radius rr and with orbital speed vm{v_m}, the orbital speed of the satellite around moon is,

vm=GMmr{v_m} = \sqrt {\frac{{G{M_m}}}{r}} …… (4)

From the equations (3) and (4), as the mass of earth is greater than the mass of moon, the condition for the orbital velocities of the satellite around earth and moon is,

ve>vm{v_e} > {v_m} …… (5)

The time period of the orbit of the satellite around earth is,

Te=2πrve{T_e} = \frac{{2\pi r}}{{{v_e}}} …… (6)

The time period of orbit of the satellite around moon is,

Tm=2πrvm{T_m} = \frac{{2\pi r}}{{{v_m}}} …… (7)

Use the condition (5) in equations (6) and (7)

As the orbital speed around earth is greater than the orbital speed around moon, the satellite will have the time period of revolution around moon Tm{T_m} to be greater than its time period around earth Te{T_e}.

The incorrect options are,

Tm{T_m} is less than Te{T_e}

Tm{T_m} is equal to Te{T_e}

Applying the condition ve>vm{v_e} > {v_m} for the orbital velocities,

Tm{T_m} is less than Te{T_e}

The above option is incorrect because the orbital speed is inversely proportional to the time period of motion.

Tm{T_m} is equal to Te{T_e}

The above option is incorrect because the condition ve>vm{v_e} > {v_m} says that the satellite cannot have same time period of revolution moon and earth.

Hence, the correct option is

Tm{T_m} is greater than Te{T_e}

Ans: Part 1

The correct option is, at a distance that is greater than r{\bf{r}}.

Part 2

The correct option is (C) Tm{{\bf{T}}_{\bf{m}}} is greater than Te{{\bf{T}}_{\bf{e}}}